# Beyond Fully Supervised Pixel Annotations: Scribble-Driven Weakly-Supervised Framework for Image Manipulation Localization

### v1

@@ Abstract
Deep learning-based image manipulation localization (IML) methods have achieved remarkable performance in recent years, but typically rely on large-scale pixel-level annotated datasets. To address the challenge of acquiring high-quality annotations, some recent weakly supervised methods utilize image-level labels to segment manipulated regions. However, the performance is still limited due to insufficient supervision signals. In this study, we explore a form of weak supervision that improves the annotation efficiency and detection performance, namely scribble annotation supervision. We re-annotated mainstream IML datasets with scribble labels and propose the first scribble-based IML (Sc-IML) dataset. Additionally, we propose the first scribble-based weakly supervised IML framework. Specifically, we employ self-supervised training with a structural consistency loss to encourage the model to produce consistent predictions under multi-scale and augmented inputs. In addition, we propose a prior-aware feature modulation module (PFMM) that adaptively integrates prior information from both manipulated and authentic regions for dynamic feature adjustment, further enhancing feature discriminability and prediction consistency in complex scenes. We also propose a gated adaptive fusion module (GAFM) that utilizes gating mechanisms to regulate information flow during feature fusion, guiding the model toward emphasizing potential tampered regions. Finally, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$). This loss dynamically regularizes predictions in weakly annotated or unlabeled regions based on model uncertainty, effectively suppressing unreliable predictions. Experimental results show that our method outperforms existing fully supervised approaches in terms of average performance both in-distribution and out-of-distribution. Our code is available on https://github.com/vpsg-research/SCAF.
Index Terms:
Image manipulation localization, weakly supervised, scribble, consistency, uncertainty.
publicationid: pubid:
I Introduction
@@
Malicious manipulated images will spread false information and seriously endanger social harmony. Thus, image manipulation localization (IML) technology, which aims at accurately segmenting the manipulated region, has been widely studied.
@@
Figure 1: Compared with the existing IML methods in labeling efficiency and performance.
@@
In recent years, deep learning techniques have significantly advanced the performance of IML methods, particularly when supported by large-scale pixel-level annotated datasets. Fully supervised approaches can leverage dense pixel annotations to learn fine-grained manipulation features, achieving state-of-the-art localization results on various benchmark datasets. However, acquiring high-quality pixel-level annotations is both time-consuming and labor-intensive, which greatly limits the scalability and practical application of these methods. In real-world scenarios, obtaining large-scale high-precision pixel annotations is often infeasible, especially as data diversity and scale continue to expand. To reduce annotation costs, weakly supervised IML methods have emerged in recent years, typically relying on image-level labels to guide the localization of manipulated regions. While such methods alleviate the dependence on dense annotations, their limited supervision signals lead to clear bottlenecks in localization accuracy and generalization capability. A substantial performance gap still exists between weakly supervised and fully supervised approaches.
@@
To address these challenges, this paper explores a form of weak supervision, namely scribble annotations, which takes into account the efficiency of labeling and the amount of supervision information, as shown in Fig. 1. Scribble annotations enable rapid labeling of large-scale IML datasets while providing informative cues for localization tasks. We re-annotated several mainstream IML datasets, including 5,123 images from CASIAv2 [1], 70 images from Coverage [2], 130 images from Columbia [3], and 414 images from NIST16 [4], resulting in a total of 5,737 images. This constitutes the first scribble-based IML (Sc-IML) dataset. During annotation, annotators used CVAT [^1] https://app.cvat.ai/ to draw scribbles on the manipulated regions based on their first impression, without referring to the ground truth. To ensure high-quality data, the annotation process was cross-verified by three reviewers. Each image took approximately 20 seconds to annotate. However, existing studies have not reported the time required to manually annotate a pixel-level mask for a manipulated image. To address this issue, we organized 10 experienced computer vision researchers to follow the same annotation process as the scribble annotations. Each researcher randomly selected 10 images from the CASIAv2 dataset for pixel-level annotation, with an average annotation time of approximately 23 minutes per image. This shows that scribble annotation is 69 times faster than pixel-level annotation. Although scribble annotations are significantly more efficient and convenient compared to pixel-level masks, they still face two major limitations: (1) Scribble annotation is a highly subjective form of weak supervision. Different annotators may have varying interpretations of the manipulated regions, boundaries, and details, resulting in significant inconsistency among the scribble annotations. (2) Scribble labels offer limited pixel-level supervision, causing the model to lack confidence in classifying unmarked regions and leading to prediction uncertainty.
@@
Based on this, we propose the first weakly supervised IML framework utilizing scribble annotations, namely SCAF. To address the challenges of prediction inconsistency and uncertainty under scribble-based weak supervision in IML datasets, we introduce a series of innovative strategies. To resolve inconsistency, we employ self-supervised training with a structural consistency loss, which constrains the model to produce consistent predictions under multi-scale and augmented inputs. We also propose a prior-aware feature modulation module (PFMM) that adaptively integrates prior information from both manipulated and authentic regions for dynamic feature adjustment, and incorporate coordinate attention to efficiently model spatial dependencies, thereby significantly enhancing the discriminability and scene adaptability of feature representations. To tackle prediction uncertainty, we propose a gated adaptive fusion module (GAFM) that regulates information flow through multi-branch channel splitting and adaptive dynamic fusion. This module not only highlights critical features and suppresses redundancy, but also leverages scribble supervision to guide the model’s attention toward potential manipulated regions. Additionally, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$), which dynamically filters model outputs based on uncertainty and applies adaptive entropy regularization to weakly annotated or unlabeled regions. This effectively suppresses unreliable predictions and improves model confidence and generalization in key areas.
In summary, our contributions to IML are as follows:
@@
  • We propose a PFMM that adaptively integrates prior information from both manipulated and authentic regions to achieve dynamic feature modulation. By further incorporating coordinate attention to efficiently model spatial dependencies, thereby significantly enhancing feature discrimination and scene adaptability.
@@
II Related Work
II-A Fully Supervised Image Manipulation Localization
With the continuous development of deep learning and the emergence of large-scale datasets, fully supervised IML methods have achieved remarkable results. For example, Guillaro et al. [5] proposed an IML framework that combines RGB data with noise-resistant fingerprints for IML. Chen et al. [6] effectively enhanced the detection of boundary artifacts in IML by refining edge features and leveraging multi-feature fusion. Kong et al. [7] leveraged pixel inconsistency to model global and local dependencies, enabling generalizable localization across various manipulation types.
II-B Weakly Supervised Image Manipulation Localization
To reduce annotation costs, weakly supervised IML methods have been developed to segment manipulated regions using only image-level labels. For example, Zhai et al. [8] introduced a self-consistency learning approach with image-level labels and multi-source consistency for improved IML. Zhou et al. [9] proposed a self-optimization framework that iteratively refines pseudolabels to enhance boundary accuracy. Li et al. [10] designed a cross-contrastive learning network that utilizes image-level labels and multi-stream feature fusion for IML. Bai [11] proposed a WSCCL 0that uses point correspondences to detect and localize manipulated regions. Notably, these are the only published studies on weakly supervised IML to date.
Although fully supervised methods achieve high localization accuracy, they rely on large-scale high-quality pixel-level annotations, which are costly and time-consuming to obtain. Their scalability is further limited by the lack of such data in real-world scenarios. Weakly-supervised IML methods using image-level labels greatly reduce annotation costs but lack spatial cues, making precise localization difficult and often missing subtle manipulations. To address these challenges, we propose a novel weakly supervised IML approach with scribble annotations. Scribbles provide crucial spatial information to the model, guiding it to learn the spatial distribution of manipulated regions while significantly improving annotation efficiency. Our method outperforms existing fully supervised methods in terms of average performance, both in-distribution and out-of-distribution.
@@ III-A Overview
The overall architecture of SCAF, is shown in Fig. 2. Specifically, the input image $\bm{I}$ is first processed by PVTv2 [12] to extract multi-scale features $\bm{f}_{i}$, and these features are then modulated in the prior-aware feature modulation module (PFMM) using the prior to integrate them into $\bm{x}_{i}$. Finally, the features are fused by the gated adaptive fusion module (GAFM) to produce the final predicted mask. The training process of the model is illustrated in Fig. 3. An image $\bm{I}$ is input into the SCAF to generate the primary output $\bm{M}_{1}$ as well as auxiliary outputs $\bm{M}_{2}$ and $\bm{M}_{3}$. Additionally, a randomly transformed version of $\bm{I}$, denoted as $T(\bm{I})$ (which may include rotations, scaling, or flipping), is also fed into the SCAF to obtain $\bm{M}_{1}^{\prime}$, which, together with $\bm{M}1$, is used to compute the structural consistency loss (${\mathcal{L}}_{{SC}}$[13]. Furthermore, $\bm{M}_{1}$ is jointly supervised by the partial cross-entropy loss (${\mathcal{L}}_{{PCE}}$), context affinity loss (${\mathcal{L}}_{{CA}}$[14], and confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$), while $\bm{M}_{2}$ and $\bm{M}_{3}$ are optimized using only ${\mathcal{L}}_{{CA}}$ and ${\mathcal{L}}_{{PCE}}$.
III-B Prior-aware Feature Modulation Module
@@
For scribble annotations, annotators typically mark the tampered regions in an image by freely drawing rough scribbles. Since this annotation method heavily relies on personal subjective judgment, different annotators may have varying understandings of the tampered regions, interpretations of boundaries, and levels of attention to detail. Even for the same image, the content of the scribble annotations can differ significantly. This subjectivity and randomness lead to considerable label inconsistency within the training set, which in turn affects the model’s ability to accurately identify tampered regions and generalize to new data. In complex scenarios, this often results in unstable and unreliable predictions. To address the inconsistency caused by subjective scribble annotations, we propose a prior-aware feature modulation module (PFMM). As shown in the lower left of Fig. 2, PFMM is composed of two parts: the manipulated discriminator (MD) and the feature modulation module (FMM). The core idea is to utilize manipulated priors (MP) and authentic prior (AP) generated by the MD, and then leverage these priors within the FMM to dynamically modulate the features $\bm{f}_{i}$ in a targeted manner. By explicitly introducing prior knowledge, PFMM can effectively mitigate the noise and bias introduced by subjective scribble annotations, guiding the model to focus on more objective region distributions. This mechanism enhances the model’s discriminability for manipulated regions and improves prediction consistency and robustness.
@@
1) Manipulated discriminator (MD): Existing research [15] has demonstrated that patch-level features are effective for detecting subtle anomalies. However, they lack explicit suppression of semantic bias, which becomes particularly evident in cross-domain scenarios. To address this limitation, we propose a selective suppression mechanism based on semantic-level features, which builds upon patch-level representations. By memorizing multi-scale semantics from both authentic and manipulated images, our method adaptively suppresses features that are highly similar to those stored in the memory bank during inference. This enables the generation of precise authentic region priors and manipulated region priors. Specifically, taking the training process of authentic images as an example, we first extract multi-layer features using the PVTv2. Each layer’s feature map is then partitioned into overlapping patches with a stride of 1, resulting in a set of local patch features ${\bm{p}_{j}^{s}}$. To further incorporate local contextual information, we perform weighted fusion of each patch and its eight spatial neighbors to obtain $\tilde{\bm{p}}_{j}$, as formulated below:
@@
The model constructs a memory bank of authentic region features by learning from a large number of real images. However, this bank is filled with highly similar features, resulting in redundancy and bias, which leads the model to overgeneralize dominant authentic region features. During inference, this mechanism causes subtle manipulated features to be easily overwhelmed by the abundance of highly matching authentic features, thereby reducing the saliency of tampered regions and ultimately lowering localization accuracy. To address this issue, we actively identify and suppress redundant authentic feature noise, which relatively amplifies and highlights manipulated features that represent inconsistencies. In this way, the model shifts from passively learning authentic image features to actively capturing the distinctive traces of manipulation. The details are as follows:
@@
where $\bm{q}$ denotes the current input feature, which is compared with each feature $\bm{v}_{m}$ in $\mathcal{B}_{s}$ to compute the maximum cosine similarity. This process suppresses regions that are overly similar to authentic features from the training set and highlights manipulated features, resulting in $\bm{q}{{sup}}$. Finally, the Euclidean distance between $\bm{q}_{sup}$ and its nearest feature $\bm{P}_{s}$ in $\mathcal{B}_{p}$ is calculated. It can be formulated as follows:
@@
where $\epsilon$ is a small constant added to prevent division by zero. This controllable probabilistic modeling approach allows the module to adaptively adjust its sensitivity to different regions based on the data distribution, effectively reducing the impact of label inconsistency or prior noise on performance. Then, $\bm{G}$ is passed through a $1\times 1$ convolution, batch normalization, and a Sigmoid activation function to generate an enhanced feature map $\bm{G}_{e}$ for each spatial location. This is further regularized residually by a learnable parameter $\gamma$ to obtain the feature $\bm{F}$, as follows:
@@
[…]
@@
III-C Gated Adaptive Fusion Module
@@
[…]
@@
Subsequently, the extracted local-detail feature $\bm{L}$ and contextual semantic feature $\bm{C}$ are concatenated along the channel dimension and fed into a gating unit $g(\cdot)$ [16]. By effectively leveraging the complementary relationship between these two types of features, the gating unit adaptively generates a more accurate gating coefficient $\alpha$, enabling precise modulation and enhancement of features of the manipulated region. It can be formulated as:
@@
Due to the limited pixel-level supervision provided by scribble-based annotations, the model may confuse features of manipulated regions with those of the authentic background, leading to ambiguous localization. To address this, we construct a complementary reverse mask ($1-\bm{O}$) from the feature map $\bm{O}$ and combine it with feature $\bm{C}$ to obtain the authentic background features. we enhance the manipulated region’s feature representation $\bm{O}$ with feature $\bm{L}$ and applying a residual connection. Subsequently, the difference between the manipulated-enhanced feature $\bm{Me}$ and the authentic background features is computed to highlight subtle or latent anomalies within the manipulated regions, resulting in the differential feature $\bm{D}$. Then, the manipulated-enhanced feature $\bm{Me}$ is fused with the differential feature $\bm{D}$ to obtain the output feature $\bm{ey}_{j}$, which can be expressed as $\bm{ey}_{k}=\operatorname{Conv1}(\bm{Me}+\bm{D})$.
@@
III-D Confidence-aware Entropy Minimization Loss
In scribble-based weakly supervised IML, the sparsity of supervision signals makes it difficult to adequately constrain the model’s learning. While entropy minimization is an effective strategy for leveraging unlabeled regions, blindly applying it may cause the model to become “over-confident” on unreliable predictions, amplifying errors and resulting in unstable training. To address this issue, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$). This loss applies entropy minimization only to unlabeled pixels whose current predictions are deemed reliable based on a confidence threshold. Additionally, a weak regularization term is applied to the annotated regions to directly encourage the model to produce predictions with higher confidence and sharper boundaries. Specifically, the model primary outputs $\bm{M_{1}}\in\mathbb{R}^{B\times 1\times H\times W}$, where each pixel produces a probability value $\bm{m}_{t}\in[0,1]$. The Shannon entropy $\mathcal{H}(\cdot)$ for each pixel is defined as follows:
@@ III-D Confidence-aware Entropy Minimization Loss
TABLE I: The dataset used in our experiments. Copy-move, splicing, and inpainting are abbreviated as CM, SP, and IP, respectively.
Dataset Nums #CM #SP #IP Train Test
CASIAv2 5123 3295 1828 0 5123 0
CASIAv1 920 459 461 0 0 920
Coverage 100 100 0 0 70 30
Columbia 180 0 180 0 130 50
NIST16 564 68 288 208 414 150
CocoGlide 512 - - - 0 512
In-the-wild 201 0 201 0 0 201
Korus 220 - - - 0 220
DSO 100 0 100 0 0 100
IMD2020 2010 - - - 0 2010
@@
where $T$ denotes the current epoch and $ramp$ represents the ramp-up period. During this period, $\lambda(T)$ increases smoothly from 0 to its maximum value $w_{max}$. In this work, we set $w_{max}$ and $w_{weak}$ to 0.1, which is chosen to balance training stability and final model performance. Thus, the final loss term is defined as $\mathcal{L}_{CEM}=\lambda(T)\cdot\mathcal{L}_{ent}$.
@@
TABLE II: Comparison with other advanced weakly supervised IML methods.
Method Pub. CASIAv1.0 Coverage NIST16 IMD2020 Columbia
WSCL ICCV’23 0.153 0.201 0.099 0.173 0.362
EdgeCAM ESWA’24 0.301 0.262 0.254 0.242 0.470
SOWCL ICASSP’25 0.334 0.239 0.288 0.259 0.385
WSCCL KBS’25 0.349 0.281 0.278 0.259 0.516
Ours - 0.716 0.827 0.721 0.580 0.979
IV Experiments and Results
IV-A Datasets
Our experiments primarily utilize 9 mainstream benchmark datasets: NIST16 [4], CASIA [1], Columbia [3], Coverage [2], IMD2020 [17], CocoGlide [18], In-the-Wild [19], DSO [20], Korus [21]. The dataset partitioning is presented in Table I.
IV-B Implementation Details

### v2

@@ Abstract
Deep learning-based image manipulation localization (IML) methods have achieved remarkable performance in recent years, but typically rely on large-scale pixel-level annotated datasets. To address the challenge of acquiring high-quality annotations, some recent weakly supervised methods utilize image-level labels to segment manipulated regions. However, the performance is still limited due to insufficient supervision signals. In this study, we explore a form of weak supervision that improves the annotation efficiency and detection performance, namely scribble annotation supervision. We re-annotate mainstream IML datasets with scribble labels and propose the first scribble-based IML (Sc-IML) dataset. Additionally, we propose the first scribble-based weakly supervised IML framework. Specifically, we employ self-supervised training with a structural consistency loss to encourage the model to produce consistent predictions under multi-scale and augmented inputs. In addition, we propose a prior-aware feature modulation module (PFMM) that adaptively integrates prior information from both manipulated and authentic regions for dynamic feature adjustment, further enhancing feature discriminability and prediction consistency in complex scenes. We also propose a gated adaptive fusion module (GAFM) that utilizes gating mechanisms to regulate information flow during feature fusion, guiding the model toward emphasizing potential manipulated regions. Finally, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$). This loss dynamically regularizes predictions in weakly annotated or unlabeled regions based on model uncertainty, effectively suppressing unreliable predictions. Experimental results show that our method outperforms existing fully supervised approaches in terms of average performance both in-distribution and out-of-distribution.
Codehttps://github.com/vpsg-research/SCAF
Introduction
@@
Maliciously manipulated images can spread false information and seriously threaten social harmony. To safeguard information security and uncover the truth behind image manipulation, image manipulation localization (IML) technology, which focuses on accurately segmenting the manipulated regions, has been extensively studied.
@@
Figure 1: Comparative results between our model and existing IML methods in labeling efficiency and performance.
@@
In recent years, deep learning techniques have significantly advanced the performance of IML methods, particularly when supported by large-scale pixel-level annotated datasets. Fully supervised methods can leverage dense pixel annotations to learn fine-grained manipulation features, achieving state-of-the-art localization results on various benchmark datasets. However, acquiring high-quality pixel-level annotations is both time-consuming and labor-intensive, which greatly limits the scalability and practical application of these methods. In real-world scenarios, obtaining large-scale high-precision pixel annotations is often infeasible, especially as data diversity and scale continue to expand. To reduce annotation costs, weakly supervised IML methods have emerged in recent years, typically relying on image-level labels to guide the localization of manipulated regions. While such methods alleviate the dependence on dense annotations, their limited supervision signals lead to clear bottlenecks in localization accuracy and generalization capability. A substantial performance gap still exists between weakly supervised and fully supervised methods.
@@
To address these challenges, this paper explores a form of weak supervision, namely scribble annotations, which takes into account the efficiency of labeling and the amount of supervision information, as shown in Fig. 1. Scribble annotations enable rapid labeling of large-scale IML datasets while providing informative cues for localization tasks. We re-annotated several mainstream IML datasets, including 5,123 images from CASIAv2 (dong2013casia), 70 images from Coverage (wen2016coverage), 130 images from Columbia (hsu2006columbia), and 414 images from NIST16 (guan2019mfc), resulting in a total of 5,737 images. This constitutes the first scribble-based IML (Sc-IML) dataset. During annotation, annotators used CVAT to draw scribbles on the manipulated regions based on their first impression, without referring to the ground truth. To ensure high-quality data, the annotation process was cross-verified by three reviewers. Each image took approximately 20 seconds to annotate. However, existing studies have not reported the time required to manually annotate a pixel-level mask for a manipulated image. To address this issue, we organized 10 experienced computer vision researchers. Each researcher randomly selected 10 images from the training dataset for pixel-level annotation, with an average annotation time of approximately 23 minutes per image. This shows that scribble annotation is 69 times faster than pixel-level annotation. Although scribble annotations are significantly more efficient and convenient compared to pixel-level masks, they still face two major limitations: (1) Scribble annotation is a highly subjective form of weak supervision. Different annotators may have varying interpretations of the manipulated regions, boundaries, and details, resulting in significant inconsistency among the scribble annotations. (2) Scribble labels offer limited pixel-level supervision, causing the model to lack confidence in classifying unmarked regions and leading to prediction uncertainty.
@@
Based on this, we propose the first weakly supervised IML framework utilizing scribble annotations, namely SCAF. To address the challenges of prediction inconsistency and uncertainty under scribble-based weak supervision in IML datasets, we introduce a series of innovative strategies. To resolve inconsistency, we employ self-supervised training with a structural consistency loss, which constrains the model to produce consistent predictions under multi-scale and augmented inputs. We also propose a prior-aware feature modulation module (PFMM) that adaptively integrates prior information from both manipulated and authentic regions for dynamic feature adjustment, and incorporate coordinate attention to efficiently model spatial dependencies, thereby significantly enhancing the discriminability and scene adaptability of feature representations. To tackle prediction uncertainty, we propose a gated adaptive fusion module (GAFM) that regulates information flow through multi-branch channel splitting and adaptive dynamic fusion. This module not only highlights critical features and suppresses redundancy, but also leverages scribble supervision to guide the model’s attention toward potential manipulated regions. Additionally, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$), which dynamically filters model outputs based on uncertainty and applies adaptive entropy regularization to weakly annotated or unlabeled regions. This effectively suppresses unreliable predictions and improves model confidence and generalization in key areas. In summary, our contributions to IML are as follows:
@@
  • We propose a PFMM that adaptively integrates prior information from both manipulated and authentic regions to achieve dynamic feature modulation. We further incorporate coordinate attention to efficiently model spatial dependencies, thereby significantly enhancing feature discrimination and scene adaptability.
@@
Related Work
Fully supervised image manipulation localization: With the continuous development of deep learning and the emergence of large-scale datasets, fully supervised IML methods have achieved remarkable results. For example, (guillaro2023trufor) proposed an IML framework that combines RGB data with noise-resistant fingerprints for IML. (chen2024ean) effectively enhances the detection of boundary artifacts in IML by refining edge features and leveraging multi-feature fusion. (10.1504/ijaacs.2024.139383) employs a Siamese network to extract illumination features and integrate them with a U-shaped network to localize manipulated regions. (10.1504/ijaacs.2024.142523) embeds manipulation cues into similarity features to guide the matching process toward suspicious regions. (10883001) exploits pixel inconsistencies to model global and local dependencies, enabling robust generalization to diverse manipulations.
Weakly supervised image manipulation localization: To reduce annotation costs, weakly supervised IML methods have been developed to segment manipulated regions using only image-level labels during training. (zhai2023towards) applies self-consistency learning with multi-source cues to improve localization. (ZHOU2024123501) iteratively refines pseudolabels to sharpen boundaries. (10889843) introduces a cross-contrastive multi-stream fusion network that dispenses with pixel-level annotations.  (bai2025weakly) proposes WSCCL, which exploits multiple point correspondences for IML.
Although fully supervised methods achieve high localization accuracy, they rely on large-scale high-quality pixel-level annotations, which are costly and time-consuming to obtain. Their scalability is further limited by the lack of such data in real-world scenarios. Weakly supervised IML methods using image-level labels greatly reduce annotation costs but lack spatial cues, making precise localization difficult and often missing subtle manipulations. To address these challenges, we propose a novel weakly supervised IML approach with scribble annotations. Scribbles provide crucial spatial information to the model, guiding it to learn the spatial distribution of manipulated regions while significantly improving annotation efficiency. Our model outperforms existing fully supervised methods in terms of average performance, both in-distribution and out-of-distribution.
@@
The overall architecture of SCAF is shown in Fig. 2. Specifically, the input image $\bm{I}$ is first processed by PVTv2 (wang2022pvt) to extract multi-scale features $\bm{f}_{i}$, and these features are then modulated in the prior-aware feature modulation module (PFMM) using the prior to integrate them into $\bm{x}_{i}$. Finally, the features are fused by the gated adaptive fusion module (GAFM) to produce the final predicted mask. The training process of the model is illustrated in Fig. 3. An image $\bm{I}$ is input into the SCAF to generate the primary output $\bm{M}_{1}$ as well as auxiliary outputs $\bm{M}_{2}$ and $\bm{M}_{3}$. Additionally, a randomly transformed version of $\bm{I}$, denoted as $T(\bm{I})$, which may include rotation, scaling, or flipping, is also fed into the SCAF to obtain $\bm{M}_{1}^{\prime}$. This, together with $\bm{M}_{1}$, is used to compute the structural consistency loss (${\mathcal{L}}_{{SC}}$(he2023weakly). Furthermore, $\bm{M}_{1}$ is jointly supervised by the partial cross-entropy loss (${\mathcal{L}}_{{PCE}}$), context affinity loss (${\mathcal{L}}_{{CA}}$(obukhov2019gated), and confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$), while $\bm{M}_{2}$ and $\bm{M}_{3}$ are optimized using only ${\mathcal{L}}_{{CA}}$ and ${\mathcal{L}}_{{PCE}}$.
Prior-aware Feature Modulation Module
@@
For scribble annotations, annotators typically mark the manipulated regions in an image by freely drawing rough scribbles. Since this annotation method heavily relies on personal subjective judgment, different annotators may have varying understandings of the manipulated regions, interpretations of boundaries, and levels of attention to detail. Even for the same image, the content of the scribble annotations can differ significantly. This subjectivity and randomness lead to considerable label inconsistency within the training set, which in turn affects the model’s ability to accurately identify manipulated regions and generalize to new data. In complex scenarios, this often results in unstable and unreliable predictions. To address the inconsistency caused by subjective scribble annotations, we propose a prior-aware feature modulation module (PFMM). As shown in the lower left of Fig. 2, PFMM is composed of two parts: the manipulated discriminator (MD) and the feature modulation module (FMM). The core idea is to utilize manipulated priors (MP) and authentic prior (AP) generated by the MD, and then leverage these priors within the FMM to dynamically modulate the features $\bm{f}_{i}$ in a targeted manner. By explicitly introducing prior knowledge, PFMM can effectively mitigate the noise and bias introduced by subjective scribble annotations, guiding the model to focus on more objective region distributions. This mechanism enhances the model’s discriminability for manipulated regions and improves prediction consistency and robustness.
@@
1) Manipulated discriminator (MD): Existing research (roth2022towards) has demonstrated that patch-level features are effective for detecting subtle anomalies. However, they lack explicit suppression of semantic bias, which becomes particularly evident in cross-domain scenarios. To address this limitation, we propose a selective suppression mechanism based on semantic-level features, which builds upon patch-level representations. By memorizing multi-scale semantics from both authentic and manipulated images, our method adaptively suppresses features that are highly similar to those stored in the memory bank during inference. This enables the generation of precise authentic region priors and manipulated region priors. Specifically, taking the training process of authentic images as an example, we first extract multi-layer features using the PVTv2. Each layer’s feature map is then partitioned into overlapping patches with a stride of 1, resulting in a set of local patch features $\bm{p}_{j}^{s}$}. To further incorporate local contextual information, we perform weighted fusion of each patch and its eight spatial neighbors to obtain $\tilde{\bm{p}}_{j}$, as formulated below:
@@
The model constructs a memory bank of authentic region features by learning from a large number of real images. However, this bank is filled with highly similar features, resulting in redundancy and bias, which leads the model to overgeneralize dominant authentic region features. During inference, this mechanism causes subtle manipulated features to be easily overwhelmed by the abundance of highly matching authentic features, thereby reducing the saliency of manipulated regions and ultimately lowering localization accuracy. To address this issue, we actively identify and suppress redundant authentic feature noise, which relatively amplifies and highlights manipulated features that represent inconsistencies. In this way, the model shifts from passively learning authentic image features to actively capturing the distinctive traces of manipulation. The details are as follows:
@@
where $\bm{q}$ denotes the current input feature, which is compared with each feature $\bm{v}_{m}$ in $\mathcal{B}_{s}$ to compute the maximum cosine similarity. This process suppresses regions that are overly similar to authentic features from the training set and highlights manipulated features, resulting in $\bm{q}_{{sup}}$. Finally, the Euclidean distance between $\bm{q}_{sup}$ and its nearest feature $\bm{P}_{s}$ in $\mathcal{B}_{p}$ is calculated. It can be formulated as follows:
@@
where $\epsilon$ is a small constant added to prevent division by zero. This controllable probabilistic modeling approach allows the module to adaptively adjust its sensitivity to different regions based on the data distribution, effectively reducing the impact of label inconsistency or prior noise on performance. Then, $\bm{G}$ is passed through a $1\times 1$ convolution, batch normalization, and a sigmoid activation function to generate an enhanced feature map $\bm{Ge}$ for each spatial location. This is further regularized residually by a learnable parameter $\gamma$ to obtain the feature $\bm{F}$, as follows:
@@
where $\operatorname{Conv1}$ denotes a $1\times 1$ convolution. Finally, coordinate attention is applied to the concatenated features of $\bm{F}$ and $\bm{f}_{i}$, resulting in the output feature $\bm{x}_{i}$. Coordinate attention captures both spatial location information and channel interdependencies, further enhancing the model’s ability to localize and discriminate manipulated regions.
@@
Gated Adaptive Fusion Module
@@
where $\operatorname{Conv3}$ denotes a $3\times 3$ convolution. ${Cat}\left.\right|_{z=k}^{3}$ denotes the concatenation of all $\bm{ex}_{n}$ for $n=j$ to 3 in sequence. $\mathbb{I}(k>0)$ denotes the indicator function, which returns 1 if $k>0$, and 0 otherwise. The GMM employs a gating mechanism to achieve adaptive recalibration and interaction between local and global features, thereby enhancing the modeling capability for fine-grained representations. Specifically, the feature $\bm{y}_{k}$ is first processed through two separate sequences of $3\times 3$ convolutional layers followed by ReLU activations, resulting in two distinct feature branches.
@@
Subsequently, the extracted local-detail feature $\bm{L}$ and contextual semantic feature $\bm{C}$ are concatenated along the channel dimension and fed into a gating unit $g(\cdot)$ (10038722). By effectively leveraging the complementary relationship between these two types of features, the gating unit adaptively generates a more accurate gating coefficient $\alpha$, enabling precise modulation and enhancement of features of the manipulated region. It can be formulated as:
@@
Due to the limited pixel-level supervision provided by scribble-based annotations, the model may confuse features of manipulated regions with those of the authentic background, leading to ambiguous localization. To address this, we construct a complementary reverse mask ($1-\bm{O}$) from the feature map $\bm{O}$ and combine it with feature $\bm{C}$ to obtain the authentic background features. We enhance the manipulated region’s feature representation $\bm{O}$ by combining it with feature $\bm{L}$ and applying a residual connection. Subsequently, the difference between the manipulated-enhanced feature $\bm{Me}$ and the authentic background features is computed to highlight subtle or latent anomalies within the manipulated regions, resulting in the differential feature $\bm{D}$. Then, the manipulated-enhanced feature $\bm{Me}$ is fused with the differential feature $\bm{D}$ to obtain the output feature $\bm{ey}_{j}$, which can be expressed as $\bm{ey}_{k}=\operatorname{Conv1}(\bm{Me}+\bm{D})$.
@@
Confidence-aware Entropy Minimization Loss
In scribble-based weakly supervised IML, the sparsity of supervision signals makes it difficult to adequately constrain the model’s learning. While entropy minimization is an effective strategy for leveraging unlabeled regions, blindly applying it may cause the model to become “over-confident” on unreliable predictions, amplifying errors and resulting in unstable training. To address this issue, we propose a confidence-aware entropy minimization loss (${\mathcal{L}}_{{CEM}}$). This loss applies entropy minimization only to unlabeled pixels whose current predictions are deemed reliable based on a confidence threshold. Additionally, a weak regularization term is applied to the annotated regions to directly encourage the model to produce predictions with higher confidence and sharper boundaries. Specifically, the model’s primary output is $\bm{M}_{1}\in\mathbb{R}^{B\times 1\times H\times W}$, where each pixel is associated with a probability value $m_{t}\in[0,1]$. The Shannon entropy $\mathcal{H}(\cdot)$ for each pixel is defined as follows:
@@
where $T$ denotes the current epoch and rampramp represents the ramp-up period. During this period, $\lambda(T)$ increases smoothly from 0 to its maximum value $w_{max}$. In this work, we set $w_{max}$ and $w_{weak}$ to 0.1, which is chosen to balance training stability and final model performance. Thus, the final loss term is defined as $\mathcal{L}_{CEM}=\lambda(T)\cdot\mathcal{L}_{ent}$.
Method CASIAv1 Coverage NIST IMD Columbia
WSCL 0.153 0.201 0.099 0.173 0.362
EdgeCAM 0.301 0.262 0.254 0.242 0.470
SOWCL 0.334 0.239 0.288 0.259 0.385
WSCCL 0.349 0.281 0.278 0.259 0.516
Ours 0.716 0.827 0.721 0.580 0.979
@@
Table 1: Comparison with other advanced weakly supervised IML methods.
Experiments and Results
Datasets and Implementation Details